Factor the following polynomial.
3x(*SQUARED) - 15x + 18
couldnt make the 2 smaller so wrote it out
should be written as "three x squared minus fifteen x plus eighteen

Answer by checkley75(3666)

3X^2-15X+18=0
(3x-6)(x-3)=0
3x-6=0
3x=6
x=2 answer.
x-3=0
x=3 answer.